OpenStudy (anonymous):

One solution of the equation 6x2 - 19x + 3 = 0 is 1/6 . What is the other solution?

OpenStudy (mertsj):

If one answer is 1/6 that means that one factor is 6x-1. Find the other factor and use it to find the other root.

OpenStudy (mertsj):

(6x-1)( )=6x^2-19x+3

jimthompson5910 (jim_thompson5910):

In ax^2+bx+c = 0, the solutions/roots p and q form this relationship p*q = c/a and p+q = -b/a We know that one root is 1/6, so p = 1/6 p*q = c/a (1/6)*q = 3/6 q/6 = 3/6 Solve this equation for q to find the other root.

OpenStudy (mikezack123):

would the answer be q = 3 ? @jim_thompson5910

OpenStudy (mikezack123):

@jim_thompson5910 ?

jimthompson5910 (jim_thompson5910):

yeah q = 3

OpenStudy (mikezack123):

thank you!

jimthompson5910 (jim_thompson5910):

np